OpenStudy (anonymous):

I am having a hard time with this question. I've ran through it several times and my answer check never pans out. Any help would be appreciated! Thanks! 2 p+1 --- + --- = 5 5p p

OpenStudy (baldymcgee6):

what do you need in order to add fractions together?

OpenStudy (anonymous):

A common denominator.

OpenStudy (baldymcgee6):

exactly so what will our common denom be?

OpenStudy (anonymous):

Is it 5p?

OpenStudy (baldymcgee6):

yes

OpenStudy (anonymous):

Thanks for your help! I'm going to give this a run

OpenStudy (baldymcgee6):

let me know if you get stuck!

OpenStudy (anonymous):

@mramirez686 Pay attention to the right side: 5 is 5/1

OpenStudy (anonymous):

I'm still running into a wall here. Anyone have an example to show me by chance?

OpenStudy (anonymous):

Show us what you got?

OpenStudy (anonymous):

@mramirez686 How do you get common denominator?

OpenStudy (anonymous):

\[\frac{ 2 }{ 5p } + \frac{ p+1 }{ p } = \frac{ 5 }{ 1 }\]Is what I have, so far given the input provided from everyone. Knowing that the common denominator I am going for is 5p. But I am doing so to cancle the fractions? I feel like I've been running around in circles with this and the examples in the book seem almost unrelated.

OpenStudy (baldymcgee6):

so how do we get the common denominator?

OpenStudy (anonymous):

factorizing and finding what is common between them?

OpenStudy (anonymous):

You already had the first denominator is 5p, now what do you multiply with the second fraction to get 5p?

OpenStudy (anonymous):

Hint: look at the second denominator!

OpenStudy (anonymous):

right, you would multiply 5 to get 5p

OpenStudy (anonymous):

Great, at least some move here! So what the second fraction becomes after multiply by 5?

OpenStudy (anonymous):

5p+5/5p

OpenStudy (anonymous):

excellent! Now what should you multiply to with the right side to get the denom. 5p?

OpenStudy (anonymous):

* multiply to the right side

OpenStudy (anonymous):

1/5p?

OpenStudy (anonymous):

What fraction on the right side?

OpenStudy (anonymous):

before you multiply?

OpenStudy (anonymous):

well, assuming 5 is the same as 5/1?

OpenStudy (anonymous):

Right, now multiply 5/ 1 with ...?

OpenStudy (anonymous):

5p, so = 25p on the right

OpenStudy (anonymous):

5 * 5P / 1 * 5P = ...?

OpenStudy (anonymous):

25p/5p

OpenStudy (anonymous):

Now, look back to make sure all the fractions have the common denom? If so, it's safely to add ....?

OpenStudy (anonymous):

yes, 5p+7/5p = 25p/5p

OpenStudy (anonymous):

would I multiply by 5p to cancel out the fractions? or am I missing a step here

OpenStudy (anonymous):

NO, the goal is COMMON DENOMINATOR = 5P!

OpenStudy (anonymous):

After you have common denom. 5p, cross out the denom! What left is numerator :)

OpenStudy (anonymous):

Is the question solving for p?

OpenStudy (anonymous):

I'm assuming, it says solve and verify.

OpenStudy (anonymous):

So what's in the numerators now?

OpenStudy (anonymous):

5p +7 = 25p

OpenStudy (anonymous):

Can you solve for p now?

OpenStudy (anonymous):

either way I do it, I am getting a fraction that isn't checking out when plugged back in

OpenStudy (anonymous):

7/20 = p

OpenStudy (anonymous):

Oh, so it does check out. Thank you so much! This has been an all day disaster for me :/

OpenStudy (anonymous):

i've ran through this problem probably a hundred different wrong ways. Knowing that the common denominator cancels out helps out a lot. Makes more sense.